OpenStudy (anonymous):

Maclaurin expansion series of sin^2 x

OpenStudy (anonymous):

Use an identity. Which I identity will I use?

OpenStudy (anonymous):

sin^2(x) = 1/2 - 1/2 cos(2x)

OpenStudy (anonymous):

http://www.math.com/tables/trig/identities.htm

OpenStudy (anonymous):

Would that be easier?

OpenStudy (anonymous):

yes, because now all you need to do is find taylor expansion of cos(x)

OpenStudy (anonymous):

why cos x? not cos 2x?

OpenStudy (anonymous):

you actually want cos(2x), but it is easier to find series for cos(x) then replace every x with 2x

OpenStudy (anonymous):

Would that work?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

What would f^(n) be?

OpenStudy (anonymous):

what series expansion of cos(x)

OpenStudy (anonymous):

maclaurin. a=0

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

I find it hard to get the f^n given 1/2-1/2cos 2x How do you do that?

OpenStudy (anonymous):

I would only try to find series expansion of cos(x)

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

Did you find it?

OpenStudy (anonymous):

yes. cos(0)+ -sin(0)/1 (x) + -cos(0)/2! (x)^2 1 -x^2/2! + x^4/4!

OpenStudy (anonymous):

that's cos(x)= \[\sum _{n=0}^{\infty } \frac{\left((-1)^nx^{2 n}\right)}{(2k)!}\]

OpenStudy (anonymous):

now replace x with 2x

OpenStudy (anonymous):

cos(2x)=\[\sum _{n=0}^{\infty } \frac{\left((-1)^n(2x)^{2 n}\right)}{(2n)!}\]

OpenStudy (anonymous):

1/2 - 1/2 cos (2x) replace cos(2x) that series

OpenStudy (anonymous):

1/2 - 1/2 (\[\sum _{n=0}^{\infty } \frac{\left((-1)^n(2x)^{2 n}\right)}{(2n)!}\])

OpenStudy (anonymous):

This is the maclaurin expansion? COOL! :D Thanks @imranmeah91!

OpenStudy (anonymous):

wait. is this a=0?

OpenStudy (anonymous):

yes, it is

OpenStudy (anonymous):

THAAAAAAAAAAAANKS! :)

OpenStudy (anonymous):

you are welcome